# Planar near-field measurements of specular and diffuse reflection of millimeter-wave absorbers

### v1

@@
Planar near-field measurements of specular and diffuse reflection of millimeter-wave absorbers
††journal: opticajournal††articletype: Research ArticleMitigating the far sidelobes of a wide field-of-view telescope is one of the critical issues for polarization observation of the cosmic microwave background. Since even small reflections of stray light at the millimeter-wave absorbers inside the telescope may create nonnegligible far sidelobes, we have developed a method to measure the reflectance of millimeter-wave absorbers, including diffuse reflections. By applying the planar near-field measurement method to the absorbers, we have enabled two-dimensional diffuse-reflection measurements, in addition to characterizing specular reflection. We have measured the reflectance of five samples (TK RAM Large and Small Tiles and Eccosorb AN-72, HR-10, and LS-22) at two angles of incidence in the frequency range from 70 GHz to 110 GHz. Compared with conventional horn-to-horn measurements, we obtained a consistent specular reflectance with a higher precision, less affected by standing waves. We have demonstrated that the angular response and diffuse-to-specular reflectance ratio differ among various materials. The measurements also imply that some absorbers may affect the polarization direction when reflecting the incident waves.
@@ 1 Introduction
The primordial gravitational waves due to inflation are believed to have left a characteristic polarization pattern called modes in the cosmic microwave background (CMB)[1]. The -mode polarization is expected to peak at large angular scales of several degrees or more, and the stray light contamination due to emission from the galactic plane entering the telescope is expected to be one of the largest systematic errors[2, 3, 4]. Stray light can enter the telescope through the far sidelobes. Next-generation CMB telescopes such as LiteBIRD need to know sidelobe to an accuracy of -56 dB[5].
@@
The reflectance properties of millimeter-wave absorbers in free space have been made using two horns and focusing optics, such as mirrors[9], in such a way that the attenuation of the coupling constant between the horns is measured (horn-to-horn method). Since stray light enters the absorber from various angles inside the telescope, the specular reflectance is measured at multiple angles of incidence[15, 16]. Also, by rotating one of the horn antennas around the absorbers, one-dimensional distributions of diffuse reflection have been measured[17, 18, 9]. However, if the scattered waves from the absorber are only partially coupled to the receiving horn, the measured reflectance is underestimated.
In this paper, we report a new method to measure the reflectance of millimeter-wave absorbers. Applying the planar near-field antenna measurement method[19, 20], we measure the amplitude and phase of the reflected waves from absorbers on a plane. By applying plane wave expansion to the measurements, we characterize the two-dimensional angular patterns of diffuse reflection, from which the specular reflectance is also evaluated.
@@ 2.1 Horn-to-horn method
Conventional millimeter-wave absorber reflection measurements have been made using two horns and focusing optics, such as mirrors[9] in such a way that the attenuation of the coupling constant between the horns is measured (horn-to-horn method; Fig. 1(a))[16, 9]. This method has been extended to one-dimensional diffuse reflection characterization by changing the angle of the receiving horn ( in Fig. 1(a))[17, 18]. However, if the scattered waves from the absorber are only partially coupled to the receiving horn, the measured reflectance is underestimated. It has been reported that the measured reflectance depends on the distance between the absorber and the receiving horn[17].
@@ 2.2 Planar near-field method
In order to accurately characterize absorber reflection, including diffuse reflection, we have applied the planar near-field measurement method[19, 22] to the reflection measurement of millimeter-wave absorbers. Figure 1(b) shows a schematic diagram of the proposed method. In this method, we illuminate the test absorber with a feed horn and measure the waves reflected from the test absorber by moving the receiver with the probe on a two-dimensional plane. The measured amplitude and phase distributions are decomposed into plane-wave components with different angular directions by plane wave expansion (two-dimensional Fourier transform) as[23, 24]
Here, is the complex electric field reflected by the test absorber; denote the spatial frequency in the plane, which relate the angle as and with being the wave number, wavelength and speed of light, respectively. is the azimuth angle from the -axis.
The planar near-field measurement has three advantages. First, by measuring the reflected wave of the absorber, we can evaluate the reflectance characteristics in more detail. Compared with the conventional horn-to-horn method, the information obtained from the two-dimensional distribution of reflected waves, including diffuse reflections, is much greater. Second, since there is no need to know the reflection point of the absorber, it is less susceptible to alignment adjustments in optical measurements. Third, compared with horn-to-horn measurements, measurements with a small-aperture probe enable more accurate measurements less affected by standing waves.
We evaluate the specular and total reflectance of the absorbers by integrating the measured two-dimensional angular reflectance into different integration ranges. The former is obtained by the integration of the reflected waves in the Half Width at Half Maximum (HWHM) of the beams of the feed horn,
@@
Here, the angular response is calculated for each frequency using Eq. (1).
@@ 3.1 Experimental setup
Figure 2 shows the measurement setup developed in this study. We measured the amplitude and phase of the electric field reflected by the test absorbers using a vector network analyzer (VNA: Keysight Technologies N5222B) and frequency extenders (Virginia Diodes WR10VNAX). Based on the reciprocity theorem, the coupling between the two horns is equivalently measured when the VNA transmitter and receiver are reversed[25]. For the convenience of the setup, we connected the probe and lens horns to the transmitter and receiver, respectively, in this experiment.
The measurements were performed with a 1 GHz interval over a frequency range of 70 GHz – 110 GHz. Signals from 11.6 to 18.3 GHz were output from the VNA and converted to six times higher frequencies. Then, the 70 – 110 GHz signal is output from the probe. We measured the reflected waves from the millimeter-wave absorber by moving the probe in 1.5 mm increments in a 75 mm 75 mm square area. The distance between the probe and test absorbers is set to 60 mm, considering mechanical clearance. This configuration measures the specular and diffuse reflection in . The reflected waves are received by a conical horn (aperture diameter 23.0 mm, opening angle 10.0∘) with a polymethyl pentene (TPX) lens (diameter 25.4 mm, focal length 50 mm) placed in the aperture. The distance between the lens horn and test absorbers is 50 mm.
The millimeter-wave absorber under test is placed on two rotating and one-axis stages. Since the movable range of the measurement stage is finite, in addition to the angle of incidence of the lens horn (AoI in Fig. 1(b)), the angle of the absorber relative to the measurement stage ( in Fig. 1(b)) was also rotated so that the measurement plane and the signal optical path are vertical. In addition, using a one-axis stage, we aligned the surface of the absorbers at a common plane for all the measurements. Considering the situation in telescopes, we measured the absorbers fixed on a 5 mm thick aluminum plate. The absorbers and the aluminum plate are 200 mm 200 mm in size. An acrylic frame held the absorber in place, leaving no space between it and the aluminum plate.
Table 1 summarizes the measurement parameters for this experiment. We measure the reflectance for two polarizations; the polarization direction of the lens horn is set parallel to the and axes in Fig. 1 (S- and P-pol, respectively). In addition, the polarization direction of the probe is set not only parallel but also perpendicular to that of the lens horn.
Frequency range | 70 - 110 GHz |
Frequency step | 1 GHz |
@@ 3.2 Millimeter-wave absorbers
Table 2 summarizes the test millimeter-wave absorbers measured in this experiment. Eccosorb AN-72[26], HR-10[27], and LS-22[28] are commercially available millimeter-wave absorbers by Laird. They are made of polyurethane-based foams with different structures, which are expected to have different diffuse reflection characteristics: LS-22 has a single-layer structure; AN-72 has a multi-layer structure; HR-10 has a single-layer structure and contains carbon powder. LS-22 is available in several thicknesses, and we measured a 3/8-inch one. Eccosorb AN-72 and Eccosorb HR-10 are used in CMB ground-based experiments[30, 31, 32]. TK RAM is a radar absorber developed by Thomas Keating for the terahertz and millimeter-wave bands, and there are two types of TK RAM[29]. TK RAM Small tile (abbreviated as TK-S) is a small tile of 25 mm 25 mm and is recommended for 200 – 600 GHz. TK RAM Large tile (abbreviated as TK-L) is a large tile of 100 mm 100 mm and recommended for 50 – 200 GHz. TK-L and TK-S have different-sized surface structures, 5.0 mm, and 2.5 mm, respectively.
@@ 3.3 Reference measurement with an aluminum plate
Assuming that the aluminum plate without the absorber is a perfect conductor[33], we measured it for intensity calibration.
Figure 3 shows reflection measurements of the reference aluminum plate. The upper panels show the angular response of the electric field parallel and perpendicular to the co-polarization of the S and P-polarized lens horn at 90 GHz. The angular response are calculated using Eq. (1), and is normalized by the peak value of . The HWHM of the lens horn was measured to be 4.34 deg at 90 GHz, consistent with the prediction by a Gaussian-beam approximation[34]. The cross-polarization level of the lens horn was measured to be dB at 90 GHz. The lower panel shows the profile in a cut-plane. We confirm that this setup can be used to measure for two polarizations since the patterns measured with both polarizations are consistent at the dB level over a range of .
@@ 4.2 Angular responses and reflectance of the absorbers
Figure 5 shows the angular responses of the reflected waves from the absorbers at 90 GHz for AoI = 30∘. The results are normalized by the peak value of the reference aluminum plate for each polarization. AN-72 has an almost isotropic attenuation from the reference aluminum plate pattern, while HR-10 has a spread pattern. LS-22 has similar patterns to AN-72, but the difference in S-pol and P-pol is smaller than that of AN-72. TK-S also shows a pronounced diffuse reflection pattern, indicating an upward tilt. This may be caused by the surface structure and errors in attaching TK-S to the reference aluminum plate.
Figure 6 shows the angular responses at 90 GHz for AoI = 45∘. It is expected that a small but direct signal from the transmitter to the receiver is present. Such a direct path between the lens horn and probe was also measured without placing a millimeter-wave absorber. As the incident angle increases, the direct path becomes more extensive, and stripe patterns due to interference with the direct path are observed more apparently than those of S-pol in the case of P-pol. This is because the sidelobe level of the lens horn is higher in the polarization direction. This may be improved in future measurements by using horns with small sidelobes, such as a corrugated horn.
@@
Figure 7 is the cross-sectional profiles of Fig. 5 for , and . While AN-72 and LS-22 have a constant attenuation from the reference aluminum plate pattern, the intensity of the reflected waves from HR-10 ranges between -40 and -20 dB regardless of the angle . Notably, in the angular range of 20 30∘, the level of the reflected waves from HR-10 is higher than that of the reference aluminum plate.
@@
Figure 8 compares the specular and total reflectance of the absorbers in the S- and P-polarization, calculated using Eqs. (2) and (3). Here, in Eq. (3) is set to 30 degrees. The gray line shows of the direct path from the probe to the lens horn measured without placing a millimeter-wave absorber. The effects of direct path is about 10 dB larger in P-pol than in S-pol. The specular reflectance of AN-72 is about -12 dB for S-pol and -17 dB for P-pol, which is about 5 dB different for the two polarizations. On the other hand, the specular reflectance of HR-10 is about -20 dB for both polarizations, which is smaller than that of AN-72. In general, dielectric materials have more significant reflectance for the P-pol than the S-pol for oblique incident wave[21]. The measurements of AN-72 show a consistent tendency. By contrast, the reflectance of LS-22 shows little difference between the two polarizations. The reflectance of TK-L was about -40 dB for both polarizations, the lowest among the five absorbers measured in this experiment.
@@
The specular reflectance (green line) of HR-10 for the S-polarization is about -20 dB, while it rises to about -15 dB for total reflectance (red line). This shows that HR-10 has larger diffuse reflection than other absorber, consistent with the isotropic spreading feature of angular patterns in Fig. 5. The total reflectance of TK-S is also higher than the specular reflectance by around 5 dB. For other absorbers, the two reflectances have little difference, indicating that diffuse reflection is small.

### v2

@@
Abstract
Mitigating the far sidelobes of a wide field-of-view telescope is one of the critical issues for polarization observation of the cosmic microwave background. Since even small reflections of stray light at the millimeter-wave absorbers inside the telescope may create nonnegligible far sidelobes, we have developed a method to measure the reflectance of millimeter-wave absorbers, including diffuse reflections. By applying the planar near-field measurement method to the absorbers, we have enabled two-dimensional diffuse-reflection measurements, in addition to characterizing specular reflection. We have measured the reflectance of five samples (TK RAM Large and Small Tiles and Eccosorb AN-72, HR-10, and LS-22) at two angles of incidence in the frequency range from 70 GHz to 110 GHz. Compared with conventional horn-to-horn measurements, we obtained a consistent specular reflectance with a higher precision, less affected by standing waves. We have demonstrated that the angular response and diffuse-to-specular reflectance ratio differ among various materials. The measurements also imply that some absorbers may affect the polarization direction when reflecting the incident waves.
@@ 1 Introduction
The primordial gravitational waves due to inflation are believed to have left a characteristic polarization pattern called modes in the cosmic microwave background (CMB)[1]. The -mode polarization is expected to peak at large angular scales of several degrees or larger, and the stray light contamination due to emission from the Galactic plane entering the telescope is expected to be one of the largest systematic errors[2, 3, 4]. Stray light can enter the telescope through the far sidelobes. Next-generation CMB telescopes such as LiteBIRD need to know sidelobe to an accuracy of -56 dB[5].
@@
The reflectance properties of millimeter-wave absorbers in free space have been made using two horns and focusing optics[9] in such a way that the attenuation of the optical coupling between the horns is measured (horn-to-horn method). Since stray light enters the absorber from various angles inside the telescope, the specular reflectance is measured at multiple angles of incidence[15, 16]. Also, by rotating one of the horn antennas around the absorbers, one-dimensional distributions of diffuse reflection have been measured[17, 18, 9]. However, if the scattered waves from the absorber are only partially coupled to the receiving horn, the measured reflectance is underestimated.
In this paper, we report a new method to measure the reflectance of millimeter-wave absorbers. Applying the planar near-field antenna measurement method[19, 20, 21], we measure the amplitude and phase of the reflected waves from absorbers on a plane. By applying plane wave expansion to the measurements, we characterize the two-dimensional angular responses of diffuse reflection, from which the specular reflectance is also evaluated.
@@ 2.1 Horn-to-horn method
Conventional millimeter-wave absorber reflection measurements have been made using two horn antennas and measuring the attenuation of the optical coupling between the horn antennas (horn-to-horn method; Fig. 1(a))[16, 9]. Since the optical coupling should be as high as possible for accurate measurements, the horn antennas are often used with focusing optics, such as mirrors[9] and lenses. This method has been extended to one-dimensional diffuse reflection characterization by changing the angle of the receiving horn ( in Fig. 1(a))[17, 18]. However, if the scattered waves from the absorber are only partially coupled to the receiving horn, the measured reflectance is underestimated. It has been reported that the measured reflectance depends on the distance between the absorber and the receiving horn[17].
@@ 2.2 Planar near-field method
To characterize two-dimensional specular and diffuse reflections of millimeter-wave absorbers, we propose to apply the planar near-field antenna measurement method to the reflection measurement of millimeter-wave absorbers. The planar near-field measurements [19, 20, 21] derive the angular radiation pattern of an antenna by the Fourier transformation of the complex electric field at the aperture of the antenna, based on the Fraunhofer diffraction [23] or van Cittert-Zernike theorem [24]. The complex electric field at the aperture is measured using a vector network analyzer by scanning a probe in a specific scan plane. To measure the field with wide angles, a small-aperture probe is used, such as an open-ended waveguide. The measured amplitude and phase distributions are decomposed into plane-wave components with different angular directions by two-dimensional Fourier transform as,
Here, is the complex electric field measured near the test antenna aperture; denote the spatial frequencies in the plane, which relate the angle as and with being the wave number; and correspond to the angle from the -axis (perpendicular to the – plane) and the azimuth angle from the -axis, respectively.
By applying the planar near-field measurement method to the reflection measurement of millimeter-wave absorbers, the angular response of the reflected waves from the absorbers, including diffuse reflection, can be acquired. In this method, we illuminate the test absorber with a horn antenna and measure the waves reflected from the test absorber by moving the receiver with the probe on a two-dimensional plane () (Fig. 1(b)). The measured amplitude and phase distributions are decomposed into the angular response using Eq. (1). Here, we defined the propagation direction of the specular reflected wave as the -axis and set the scan plane () perpendicular to the -axis. The -axis is set in the S-polarization direction. Note that P-polarized waves oscillate with the electric field parallel to the plane of incidence, while S-polarized waves oscillate perpendicular to the plane of incidence.
The planar near-field measurement has three advantages. First, by measuring the reflected wave of the absorber, we can evaluate the reflectance characteristics in more detail. Compared with the conventional horn-to-horn method, the information obtained from the two-dimensional distribution of reflected waves, including diffuse reflections, is much greater. Second, since there is no need to know the reflection point of the absorber, it is less susceptible to alignment adjustments in optical measurements. Third, compared with horn-to-horn measurements, near-field measurements are less affected by standing waves because of the small aperture probe with a small optical coupling.
We evaluate the specular and total reflectance of the absorbers by integrating the measured two-dimensional angular responses into different integration ranges. The former is obtained by the integration of the reflected waves in the half width at half maximum (HWHM) of the beams of the horn antenna,
@@
Here, and are the angular responses calculated from Eq. (1) for the perfect electric conductor (PEC) and test absorber, respectively. For the case of our experimental verification described in Sections 3.3 and 4.2, the HWHM and are and , respectively.
@@ 3.1 Experimental setup
Figure 2 shows the measurement setup developed in this study. We measured the amplitude and phase of the electric field reflected by the test absorbers using a vector network analyzer (VNA: Keysight Technologies N5222B) and frequency extenders (Virginia Diodes WR10VNAX)[25]. Based on the reciprocity theorem, the optical coupling between the two horns is equivalently measured when the VNA transmitter and receiver are reversed[26]. For the convenience of the setup, we connected a probe and a lens horn to the transmitter and receiver, respectively, in this experiment.
The measurements were performed with a 1 GHz interval over a frequency range of 70 GHz – 110 GHz. Signals from 11.6 to 18.3 GHz were output from the VNA and up-converted to six times higher frequencies. Then, the 70 – 110 GHz signal was output from the probe. We used an open-ended waveguide as the probe[27]. We measured the reflected waves from the millimeter-wave absorber by scanning the probe in 1.5 mm increments in a 75 mm 75 mm square area. The distance between the probe and test absorbers was set to 60 mm, considering the mechanical clearance. This configuration measures the specular and diffuse reflection in . The reflected waves were received by a lens horn, which consists of a conical horn (aperture diameter 23.0 mm, opening angle 10.0∘) with a polymethyl pentene (TPX) lens (diameter 25.4 mm, focal length 50 mm) placed in the aperture. The distance between the lens horn and test absorbers was 50 mm. Based on a Gaussian-beam approximation[28], the beam radius at the absorber is 10.7 mm.
The millimeter-wave absorber under test was placed on two rotating and one-axis stages. Since the movable range of the measurement stage was finite, in addition to the angle of incidence of the lens horn (AoI in Fig. 1(b)), the angle of the absorber relative to the measurement stage was also rotated so that the measurement plane (– plane) and the signal optical path (-axis) becomes perpendicular. In addition, using a one-axis stage, we aligned the surface of the absorbers at a common plane for all the measurements. Considering the situation in telescopes, we measured the absorbers fixed on a 5 mm thick aluminum plate. The absorbers and the aluminum plate were 200 mm 200 mm in size. An acrylic frame held the absorber in place, leaving no space between it and the aluminum plate.
Table 1 summarizes the measurement parameters for this experiment. We measured the reflectance for two polarizations; the polarization direction of the lens horn was set parallel to the and -axes in Fig. 1 (S- and P-pol, respectively). We checked the mechanical accuracy of the coordinates using a level gauge. In addition, the polarization direction of the probe was set not only parallel but also perpendicular to that of the lens horn.
Frequency range (step) | 70 - 110 GHz (1 GHz) |
Scan range (step) | 75 mm 75 mm (1.5 mm) |
@@ 3.2 Millimeter-wave absorbers
Table 2 summarizes the test millimeter-wave absorbers measured in this experiment. Eccosorb AN-72[29], HR-10[30], and LS-22[31] are commercially available millimeter-wave absorbers by Laird. They are made of polyurethane-based foams with different structures, which are expected to have different diffuse reflection characteristics: LS-22 has a single-layer structure; AN-72 has a multi-layer structure; HR-10 has a single-layer structure and contains carbon powder. LS-22 is available in several thicknesses, and we measured a 3/8-inch one. Eccosorb AN-72 and Eccosorb HR-10 are used in CMB ground-based experiments[33, 34, 35]. TK RAM is a radar absorber developed by Thomas Keating for the terahertz and millimeter-wave bands, and there are two types of TK RAM[32]. TK RAM Small tile (abbreviated as TK-S) is a small tile of 25 mm 25 mm and is recommended for 200 – 600 GHz. TK RAM Large tile (abbreviated as TK-L) is a large tile of 100 mm 100 mm and recommended for 50 – 200 GHz. TK-L and TK-S have different-sized surface structures with a pyramidal shape, their heights being 5.0 mm and 2.5 mm, respectively.
@@ 3.3 Reference measurement with an aluminum plate
For the intensity calibration, we measured the angular response of an aluminum plate without the absorber. Assuming that aluminum is a perfect electric conductor[36], the angular response of the aluminum plate is equivalent to that of the perfect electric conductor ( in Eqs. (2) and (3)). Hereafter, we use instead of .
Figure 3 shows the reflection measurements of the reference aluminum plate at 90 GHz, measured with the electric field parallel and perpendicular to the co-polarization of the S and P-polarized lens horn. The angular response are calculated using Eq. (1) and are normalized by their peak value. Here, we confirmed that the angular response of the reference aluminum measurement peaks at .
These patterns are consistent with the lens horn pattern. The lens horn has larger sidelobes in the electric-field direction of the co-polarization pattern and quadrupoles in the cross-polarization pattern. Therefore, as expected, the “ss” and “pp” patterns have sidelobes in each polarization direction, and the “sp” and “ps” patterns show four lobes. The HWHM of the lens horn was measured to be 4.3 ∘ at 90 GHz, consistent with the prediction by a Gaussian-beam approximation[28]. The cross-polarization level of the lens horn was measured to be dB at 90 GHz. Since the patterns measured with both polarizations are consistent at the dB level over a range of , we confirmed that this setup can measure two polarizations.
@@ 4.2 Angular responses and reflectance of the absorbers
Figures 5 and 6 show the angular responses of the reflected waves from the absorbers at 90 GHz for AoI = 30∘ and , respectively. The results are normalized by the peak value of the reference aluminum plate for each polarization. AN-72 has an almost isotropic attenuation from the reference aluminum plate pattern, while HR-10 has a spread pattern. LS-22 has similar patterns to AN-72, but the difference in S-pol and P-pol is smaller than that of AN-72. TK-S also shows a pronounced diffuse reflection pattern, indicating an upward tilt. This may be caused by the surface structure and errors in attaching TK-S to the reference aluminum plate.
@@
Figure 7 is the cross-sectional profiles of the measurements for AoI (Fig. 5) in the and directions. While AN-72 and LS-22 have a constant attenuation from the reference aluminum plate pattern, the intensity of the reflected waves from HR-10 ranges between -40 and -20 dB regardless of the angle . Notably, in the angular range of 20 30∘, the level of the reflected waves from HR-10 is higher than that of the reference aluminum plate.
@@
Figure 8 and Table 3 compare the specular reflectance of the absorbers in the S- and P-polarization, calculated using Eq. (2). The specular reflectance of AN-72 is -14 dB for S-pol and -19 dB for P-pol at 90 GHz, which is 5 dB different for the two polarizations. On the other hand, the specular reflectance of HR-10 is -22 dB for the S-polarization, which is smaller than that of AN-72. In general, dielectric materials have more significant reflectance for the P-pol than the S-pol for oblique incident wave[22]. The measurements of AN-72 show a consistent tendency. By contrast, the reflectance of LS-22 shows little difference between the two polarizations. The reflectance of TK-L was about -40 dB for both polarizations, the lowest among the five absorbers measured in this experiment.
@@
Figure 8 and Table 3 also show the total reflectance calculated using Eq. (3). Here, in Eq. (3) is set to 30∘. While the specular reflectance (green line in Fig. 8) of HR-10 for the S-polarization is -22 dB, it rises to -15 dB for total reflectance (red line). This shows that HR-10 has a larger diffuse reflection than other absorbers, consistent with the isotropic spreading feature of angular patterns in Fig. 5. The total reflectance of TK-S is also higher than the specular reflectance by around 5 dB. For other absorbers, the two reflectances have little difference, indicating that diffuse reflection is small.
ss | pp | |||
Samples | Specular | Total | Specular | Total |
AN-72 | -14 dB | -14 dB | -19 dB | -18 dB |
HR-10 | -22 dB | -15 dB | -21 dB | -19 dB |
LS-22 | -11 dB | -11 dB | -12 dB | -12 dB |
TK-L | -43 dB | -39 dB | -45 dB | -38 dB |
TK-S | -25 dB | -21 dB | -25 dB | -20 dB |
(direct coupling) | -65 dB | -54 dB | -56 dB | -41 dB |
4.3 Effects of direct coupling
In the conventional horn-to-horn methods, the minimum measurable reflectance is limited by the direct coupling from the transmitting horn to the receiving horn, especially for measurements with larger angles of reflection[17].
We found that the near-field method has similar limitations. The measured angular responses for AoI = 45 ∘ (Fig. 6) show a similar feature with vertical stripes at , regardless of the test absorber. This is because of interference between the reflected waves and the direct coupling from the lens horn to the probe. Such a feature is more apparent in the P-pol measurements, which is consistent with the higher sidelobe level of the lens horn in its polarization direction.
We evaluated the direct coupling in our experimental setup by performing near-field measurements without placing the millimeter-wave absorbers. The results are shown in Fig. 8 (gray lines) and in Table 3. In the case of P-pol measurements, the measured total reflectance of TK-L is close to the direct coupling, implying that the measured value may only represent an upper limit.
In future experiments, the accuracy of the measurement may be improved by reducing the direct coupling. For example, a horn with a small sidelobe such as a corrugated horn could be used. Also, the time-domain filtering in the near-field measurement[6] can be helpful to distinguish the direct coupling from the reflected waves from the absorber and remove the effects of the direct coupling.
4.4 Comparison with horn-to-horn method